OpenStudy (anonymous):

how many positive. negative and complex zeros are in this equation? x^3+17x^2+72x-90 ?

OpenStudy (anonymous):

you mean actual ones or possible ones?

OpenStudy (anonymous):

whats the difference between them? the question just says "your function has _______ positive/ negative/ complex zeros"

OpenStudy (anonymous):

oh in this case you would simply graph

OpenStudy (anonymous):

one positive, 2 complex

OpenStudy (anonymous):

how? my lesson doesnt say to graph. r

OpenStudy (anonymous):

i thought maybe you were supposed to use descartes rule of sign

OpenStudy (anonymous):

yeah i am

OpenStudy (anonymous):

ok, but descartes rule of sign tells you the POSSIBLE number of positive, negative, and complex roots

OpenStudy (anonymous):

we start with \[f(x)=x^3+17x^2+72x-90\] and see only one change in sign, from +72 to -90

OpenStudy (anonymous):

this tells you there is one postive root

OpenStudy (anonymous):

now replace x by -x to get \[f(-x)=(-x)^3+17(-x)^2+72(-x)-90\] \[f(x)=-x^3+17x^2-72x-90\] and here i count two changes in sign

OpenStudy (anonymous):

from -1 to +17 and from +17 to -72 that means either there are two negative roots, or no negative roots

OpenStudy (anonymous):

so your possibilites are 1) one postive, two negative 2) one postive, two complex

OpenStudy (anonymous):

it is a poly of degree 3 so there must be 3 (including the complex ones) and since it is of odd degree there must be one real one but other than that, you don't know unless you graph or check somehow

OpenStudy (anonymous):

if it was not clear what i meant, there is an excellent explanation here http://www.purplemath.com/modules/drofsign.htm

OpenStudy (anonymous):

i understand where it talks about the number of sign changes but it also says to subtract 2 if possible. thats what i dont understand

OpenStudy (anonymous):

if there are say 5 changes in sign for f(x) then there are 5 or 3 or 1 positive zero

OpenStudy (anonymous):

if there are 4 changes in sign for f(x) then either 4 or 2 or none. that is what it is telling you

OpenStudy (anonymous):

so it goes by odd numbers?